Suppose that “Colgate” sells its standard size toothpaste for Ksh 35. Its sales have been on an average

9000 units per month. Recently, its close competitor “Close–Up” reduced the price of its standard size

toothpaste from Ksh 45 to Ksh 40. As a result, Colgate’s sales declined by 2500 units per month.

(i) What is the own-price elasticity of Colgate when the price rises from Ksh 35 to Ksh 40?

(ii) Calculate the cross elasticity of demand between the two products on the market and show how

the two products are related.


1
Expert's answer
2022-02-07T17:11:29-0500

(i) What is the own-price elasticity of Colgate when the price rises from Ksh 35 to Ksh 40?

The percentage change in the quantity demanded of an item or service divided by the percentage change in the price is the own-price elasticity of demand. This demonstrates how responsive the quantity demanded is to price changes.

New sales 9000-2500=6,500

ΔQ = decline value = 2500

Original, Q = 9000

ΔP = 40-35 = 5

Original, P = 35


own-price elasticity = "\frac{ΔQ}{Q}×\frac{P}{ΔP}"


= "\frac{2500}{9000}×\frac{35}{5}"

= 1.94

Own-price elasticity is 1.94

(ii) Calculate the cross elasticity of demand between the two products on the market and show how the two products are related.


Cross Elasticity="\frac{ΔQ}{Q}×\frac{P}{ΔP}"

Colgate sells 9000 units of toothpaste per month when the price is Ksh.35

Competitor reduced price from ksh.45 to ksh.40

Colgate sales declined by 2500 per month

New sales 9000-2500=6,500

ΔQ = decline value = 2500

Original, Q = 9000

ΔP = 45-40 = 5

Original, P = 45

Cross elasticity = "\frac{2500}{9000}×\frac{45}{5}"

= 2.5

Show how the two products are related.

Cross elasticity of demand = 2.5

Since, 2.5 > 0; the goods are substitutes.

The two toothpaste are substitutes since the cross elasticity of demand is more than zero. Both Colgate and “close Up” toothpaste are substitutes in the market